OpenStudy (anonymous):

Can someone help me step by step to solve for a : 1/a + 1/2a = 1/a-3

OpenStudy (mrnood):

Put the terms on the left over a common denominator - you can then reduce the left to 1 term. 'Cross multiply' to put the terms with 'a' onto the numerator Expand any brackets you have. Move all a terms to one side, and numeric to the other. Job done

OpenStudy (unklerhaukus):

is it 1/a + 1/(2a) = 1/(a-3) or 1/a + (1/2)a = (1/a)-3 ?